# Towards Scalable Multi-Chip Wireless Networks with Near-Field Time Reversal

### v2

@@ VI-A Impact of Communication Frequency
To assess the degree of spatial focusing of our technique, we evaluate the ratio of the power in the desired target to the power of the rest of ports of the system. As observed in Fig. 7(c) for a particular link, we observe how TR focuses the energy on the target destination, as the power is superior to the sum of the powers in all the other ports. In contrast, the non-TR case will probably lead to large CCI, since the power at the intended destination is just a small fraction of the power arriving to all other ports. The difference between both cases is around 3 with small variation when changing the operation frequency of the wireless network. Hence, this confirms that TR could be used simultaneously to frequency multiplexing in a single package.
@@ VI-B Impact of Number of Chiplets
Fig. 8(b) shows the comparison of the power ratio for 4 and 16 chiplets in two different links. For the four chiplet case we take the channel GC as example. To perform a fair comparison with the 16 chiplet case, we find a random link that are in a close physical position to the 4 chiplet counterparts, which would be links in this case. In Fig. 3, we can see the position of antennas for the both cases.
@@ VII-A Zero-Order Hold Filter
Since the amplitude of the received multi-path components is preserved until the next point of sampling while maintaining constant energy level on the particular sampling period, the ZOH filtering has minor degradation of the focused energy concentration with reduced sampling rates. We next evaluate the impact of reducing the sampling frequency.
@@ VII-B Performance Evaluation
To evaluate the performance of the ZOH technique, we summarize the results of several evaluations in Fig. 11. First, we integrate the energy around the peak with a window of 0.4 ns (corresponding to a modulation speed of 2.5 Gb/s) as in Equation (10) and compare it to the energy outside of the peak, as an indication of how well the temporal concentration is maintained when sampling at different frequencies. As shown in Fig. 11(a), the energy is rather constant until the sampling rate drops below 350 GHz. As we shift further down, there is a degradation of the ratio, yet without a drop of more than 7 dB until 50GHz sampling frequency.
@@
Furthermore, the effect of interference due to muti-path was measured by varying the window size of the interference measurement. This result provides us with an approximate idea on measured signal strength when the data rate is increased (and then the window reduced). From the results of Fig. 11(b), it is apparent that when the window size is increased, the amount of captured energy also increases proportionally. In the figure we see that the impact of the window size is more noticeable at small sizes, since the finite sampling rates causes the peak to broaden. The impact of the resonance frequency is mild.
@@
Finally, to analyze the performance of the filter with BER, the data rates were evaluated with varied sampling frequencies (i.e. 400 GHz, 300 GHz, 200 GHz, 100 GHz, and 80 GHz) for three resonance frequencies. As illustrated in Fig. 11(c) for a fixed BER of 0.1, the achievable data rates seem to degrade gracefully as we reduce the sampling rate by significant amounts, still achieving remarkable speeds for a sampling rate of 80 GHz.
VIII Conclusion
@@
In this paper, we have presented the idea of applying time reversal to mitigate the sources of interference in static wireless communication scenarios within computing packages. To explore the effect of TR in varied setups, we evaluated the technique in a base scenario with four chiplets and antennas operating at 140 GHz, assuming idea TR filters, as well as for multiple variations in frequency, geometry, and also employing TR filters with finite sampling rate. Overall, the results proved the superior capabilities of TR in transmission, obtaining a remarkable amount of power concentration in targeted destinations. The performance of TR in transmission was assessed by calculating the error rate as a function of the data rate. We observed that TR provides an order-of-magnitude improvement with respect to non-TR transmissions, and that allows to perform 2–3 concurrent transmissions to further boost the data rate. We did not observe a clear variation with frequency, because the environment under evaluation is reverberant at all the evaluated frequencies. When increasing the number of chiplets, the changes were significant due to the higher complexity of a channel with numerous changes of propagation medium. Finally, we evaluated the impact of a finite sampling frequency in the TR filter on the performance of the technique. We observed a rather graceful degradation of the achievable data rate. In future work, we aim to explore the limits of the technique as we further reduce the sampling rate and further simplify the filter.

### v3

@@ VI-A Impact of Communication Frequency
To assess the degree of spatial focusing of our technique, we evaluate the ratio of the power in the desired target to the power of the rest of ports of the system. As observed in Fig. 8(c) for a particular link, we observe how TR focuses the energy on the target destination, as the power is superior to the sum of the powers in all the other ports. In contrast, the non-TR case will probably lead to a large CCI, since the power at the intended destination is just a small fraction of the power arriving to all other ports. The difference between both cases is around 3 with small variation when changing the operation frequency of the wireless network. Hence, this confirms that TR could be used simultaneously to frequency multiplexing in a single package.
@@ VI-B Impact of Number of Chiplets
Fig. 9(c) shows the comparison of the power ratio for 4 and 16 chiplets in two different links. For the four chiplet case we take the channel GC as example. To perform a fair comparison with the 16 chiplet case, we find a random link that are in a close physical position to the 4 chiplet counterparts, which would be links in this case. In Fig. 4, we can see the position of antennas for the both cases.
@@ VII-A Zero-Order Hold Filter
Since the amplitude of the received multipath components is preserved until the next point of sampling while maintaining constant energy level on the particular sampling period, the ZOH filtering has minor degradation of the focused energy concentration with reduced sampling rates. We next evaluate the impact of reducing the sampling frequency.
@@ VII-B Performance Evaluation
To evaluate the performance of the ZOH technique, we summarize the results of several evaluations in Fig. 12. First, we integrate the energy around the peak with a window of 0.4 ns (corresponding to a modulation speed of 2.5 Gb/s) as in Equation (10) and compare it to the energy outside of the peak, as an indication of how well the temporal concentration is maintained when sampling at different frequencies. As shown in Fig. 12(a), the energy is rather constant until the sampling rate drops below 250 GHz. As we shift further down, there is a degradation of the ratio, yet without a drop of more than 6 dB until 50 GHz sampling frequency.
@@
Furthermore, the effect of interference due to mutipath was measured by varying the window size of the CIR for the interference measurement. This result provides us with an approximate idea on measured signal strength when the data rate is increased (and then the window reduced). From the results of Fig. 12(b), it is apparent that when the window size is increased, the amount of captured energy also increases proportionally. In the figure we see that the impact of the window size is more noticeable at small sizes, since the finite sampling rates causes the peak to broaden. The impact of the resonance frequency is mild.
@@
Finally, to analyze the performance of the filter with BER, the data rates were evaluated with varied sampling frequencies (i.e. 200 GHz, 100 GHz, 80 GHz, and 50 GHz) for three resonance frequencies. As illustrated in Fig. 12(c) for a fixed BER of 0.1, the achievable data rates seem to degrade gracefully as we reduce the sampling rate by significant amounts, still achieving remarkable speeds for a sampling rate of 80 GHz.
VIII Discussion
TR filter is a novel approach for on-chip communication which simultaneously mitigate ISI and CCI, by leveraging the dispersive nature of the channel. Acting as a spatial matched filter, TR enhances the receive signal strength, while achieving near-perfect energy focusing at the intended receiver. Despite the performance analysed from ideal to non-ideal filter implementation, it is important to note that in real hardware implementation could impose limitations and this work represents the performance metrics by considering the initial exploration of TR filter in diverse environments.
Noise Figure at Transceiver
The reported CMOS receiver front-ends in this frequency band [35, 36, 37] exhibit noise figures of 5 to 15 dB, phase noise around –110 to –80 dBc/Hz at 1 MHz offset, and DC–Effective Isotropic Radiated Power (EIRP) efficiencies of 10–20%. Based on the data in [35], we have modeled the phase noise as discussed in [38] and the results are illustrated Fig. 13(c). Moreover, by considering the sampling jitter, each sample of the CIR is moved +0.5 times the sampling period (), to observe the effect of performance. The results in Fig 13(c) shows that the phase noise has more impact than the effect of jitter, as the sampling period is in nano-seconds in ideal TR filter response. However, these characteristics indicate that the real noise power should be scaled by roughly 10 dB with imperfect transceivers at 140 GHz (baseline configuration) and this work represents the first results representing the upper bound of TR filter performance.
Area and Power
In considering the area and power cost to implement the TR filter, the recent literature reports an on-chip 150 GS/s DAC [39] with an energy efficiency of 0.1656 aJ/bit. Moreover, as discussed in [40] and [41], DACs operating beyond 100 GS/s typically occupy an area in the range of 2.5-3.5 mm2. This capability provides us an open door to explore the sampling rates beyond 100 GS/s for TR filters, which gives data rates up to 25 Gb/s as discussed in Section VII.
Scalability
In this work, we analysed two chiplet configurations, (i) a 4 chiplet interposer with three antennas per each chiplet and (ii) a 16 chiplet interposer with a single antenna per each chiplet. The 4-chiplet configuration enabled analysis of inter and intra-chiplet communication, including antenna coupling, interference management through collective spectrum sharing, and link performance. The 16-chiplet configuration was used to evaluate communication performance at higher chiplet densities, focusing on dense spatial focusing enabled by TR filtering. Although these configurations differed in antenna density and chiplet count, the observed results exhibit qualitative and predictable trends from small to large scale systems [42].
Variability
Robustness is a crucial aspect in practical implementation of TR filters. The effective wavelength of the EM waves in a silicon interposer changes from 0.44-0.63 mm in the range of frequencies 140-200 GHz explored in this work. In considering geometry of our structure which is with chiplets, the dimensional and temperature variations, interconnect/bump randomness are in order of a few micro-meters, less than 0.5% of the wavelength. To access the resilience, we have simulated a percentage expansion of length in Cu layer, considering it has the highest thermal expansion coefficient [43][44] among the materials in a computing package. Considering the temperatures drift of a computer processor, the maximum expansion would be of 1-2%. Yet to confirm the robustness, we simulate 5% and 10% of expansion. Fig. 13(a) and Fig. 13(b) show the measured CIR for the different lengths and the corresponding energy concentration with the original TR filter for the 140 GHz case. It is observed that the figures show minor deviations based on the expansions with robust behavior upon variability to temperature drifts and dimension tolerance.
IX Conclusion
@@
In this paper, we have presented the idea of applying TR filters to mitigate the sources of interference in static wireless communication scenarios within computing packages. To explore the effect of TR in varied setups, we evaluated the technique in a base scenario with four chiplets and antennas operating at 140 GHz, assuming ideal TR filters, as well as for multiple variations in frequency, geometry, and also employing TR filters with finite sampling rate. We observed that TR provides an order-of-magnitude improvement with respect to non-TR transmissions, and that allows to perform 2–3 concurrent transmissions to further boost the data rate. The results showed no clear frequency dependence because the environment was reverberant across frequency sweep, while increasing the number of chiplets caused significant changes due to the greater channel complexity with numerous changes of propagation medium. In future, we aim to explore the limits of the TR filter in terms of sampling and resolution, while analyzing the impact of hardware impairments on the transceiver. Moreover, we plan to, investigate system-level integration challenges, such as (i) achieving low-power [45] and low-area designs suitable for large-scale multi-chip deployments and (ii) how the existing TRMAC protocol [25] can be co-optimized with the TR filter to improve efficiency and scalability in distributed systems.